OpenStudy (anonymous):

Find the circumference of the circle. Use 3.14 for . 30m

OpenStudy (anonymous):

hey!

OpenStudy (anonymous):

do you know what is the formula for circumference of a circle?

OpenStudy (anonymous):

Use formula for circumference c=2(3.14)r

OpenStudy (triciaal):

the diameter is twice the radius. the question is not complete. Does not state if you were given the radius or the diameter.

OpenStudy (anonymous):

the radius was 30 m

OpenStudy (anonymous):

This question got me stumped guys

OpenStudy (anonymous):

\(\huge \sf Circumference =2 \times \pi \times radius \)

OpenStudy (anonymous):

\(\huge \pi = 3.14 \)